# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

### v1

@@ 4.2 Evaluating the correlation
To determine whether LMs surprisal correlates with different metaphor novelty annotations, we measure multiple correlation metrics between surprisal and metaphor novelty scores/labels. For continuous novelty scores (as in VUA-ratings), we compute Pearson’s $r$ pearson1895regression and Spearman’s $\rho$ spearman1904association correlations. For binary labels (conventional vs. novel), we compute the Rank-biserial $r_{b}$ glass1966rankbiserial correlation. Rank-biserial estimates the probability that a random observation from the set of novel metaphors has a larger surprisal than one from the set of conventional metaphors, minus the reverse probability. We also estimate the potential of surprisal as a discriminator for binary novelty labels using the Area Under the ROC Curve (AUC) fawcett2006roc. We also compute the significance of all these estimates. All our metrics and tests do not assume normality, except for Pearson’s, for which we also provide a non-parametric alternative (Spearman). See Appendix x for the formulas of the computed measures.
@@ 4.3 Settings
We examine three families of decoder-only causal LMs (GPT-2, Llama 3 and Qwen2.5) gpt2; llama3; qwen2. For each model family, we select 3-4 different sizes, to represent the effect of model size on the correlation. To investigate the effect of instruction-tuning on the correlation, we include the instruction-tuned variants of Llama 3 and Qwen2.5. In [ADD TABLE], we list all the used model variants with important metadata.
@@
As explained in Section 3, we perform experiments on four datasets (Tables 6 and 7). As VUA-ratings has continuous novelty scores, we evaluate the correlation of its continuous scores to surprisal. In addition, we convert these scores to binary labels using a threshold of $0.5$ and evaluate their correlation to surprisal, allowing us to compare the results of VUA-ratings to VUA-dictionary, Lai2009 and GPT4o-metaphors, which only have binary novelty annotations.
@@ 4.4 Summary
Our experiments rely on a large collection of surprisal measures for each metaphor word in the four datasets: 32 different surprisal values (including direct and cloze) for 15,155 metaphoric words of the VUA dataset and for 208 and 200 metaphoric words of Lai2009 and GPT4o-metaphors, respectively. In Figure 3, we plot the distributions of surprisal values by model, in addition to metaphor novelty scores from VUA. From these plots and further normal distribution tests, we find that novelty scores and surprisal values are not strictly normally distributed.
@@ 5.1 Overall Correlation Results
Generally, we find LMs’ direct surprisal values correlate positively with metaphor novelty annotations across the four datasets (Tables 1 and 2). All correlation estimates are statistically significant for both the corpus-based and synthetic datasets  [^6] The relatively large corpus-based datasets (15,155 datapoints) can bias the significance values; however, we get similar significance values for the relatively small synthetic datasets (100 datapoints).. Overall for direct surprisal, the largest Pearson correlation $r=.419$, largest Spearman correlation $\rho=.417$, largest rank-biserial correlation $r_{b}=.638$ and largest AUC $auc=.819$ come from GPT2-base on VUA-ratings.
We also find that the correlation strengths’ scales differ across the different annotation datasets. By comparing the rank-biserial estimate $r_{b}$ across the four datasets, we find its ranges as follows: (.47-.64) for VUA-ratings, (.43-.58) for VUA-dictionary, (.28-.50) for Lai2009 and (.38-.63) for GPT4o-metaphors. These results show that surprisal correlates relatively more with the corpus-based datasets than the synthetic datasets. Also, surprisal correlates relatively more with human ratings on VUA than the dictionary-based annotation approach. While the most controlled dataset (Lai2009) gets the weakest correlations with surprisal.
@@
Table 3: Instruction-tuning % gains (over corresponding base variants) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT4o-metaphors.
@@ 5.2 Model Size Effects
The effect of model size is consistent (but diverging in direction) across the two dataset types. We plot the rank-biserial correlations of the ten model variants for the four datasets in Figure 2. The two corpus-based datasets (blue) show a monotonically decreasing correlation as model size increases per model family. On the other hand, the two synthetic datasets (red) exhibit an increasing correlation behaviour as model size increases, with a single minor exception at the GPT2 model family on the GPT4o-metaphors dataset.
@@ 5.4 Cloze-surprisal
In Table 4, we report the percentage gains in rank-biserial correlations for the cloze-surprisal values over the direct-surprisal values from the same model variants across the four datasets. We find cloze-surprisal to be boosting the correlations with extra points in many cases; however, it also deteriorates the correlations heavily in other cases.
For the GPT2 model family, cloze-surprisal consistently improves the corpus-based datasets correlations, achieving the strongest correlations in this study (out of the 32 recorded surprisals on each dataset) coming from GPT2-base on VUA-ratings with values of $r=.499$, $\rho=.499$, $r_{b}=.687$ and $auc=.843$. However, it deteriorates the correlations at large scales on the synthetic datasets.
Llama and Qwen model families also show positive gains in most of the cases. Most importantly, unlike the GPT2 model family, it boosts correlations on the synthetic datasets except for a few cases. Cloze-surprisal achieves the overall strongest correlation on the Lai2009 dataset using the Qwen-14B model, with $r_{b}=.539$ and $auc=.779$. Also, it achieves the overall strongest correlation on the GPT4o-metaphors dataset using the Llama-8B model, with $r_{b}=.684$ and $auc=.842$.
Model VUA-r VUA-d Lai GPT4o
@@
Table 4: Cloze-surprisal % gains (over direct-surprisal from the same model) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT4o-metaphors.
@@ 6 Discussion
Across multiple novelty annotation setups—continuous human ratings, dictionary-based binary labels, linguist-designed contrasts, and LLM-generated conventional vs. novel senses—we find consistently moderate associations between LM surprisal and metaphor novelty (best: $r=.49,\rho=.50,r_{b}=.69,\mathrm{AUC}=.84$). While these results are not directly comparable to prior work due to differences in datasets and task formulations, their magnitudes are broadly in line with reported surprisal–behavior correlations in acceptability judgments tjuatja-etal-2025-goes and reading times oh-schuler-2023-transformer. The closest point of reference in the metaphor novelty literature is parde_corpus_2018, who report $r=.44$ when predicting novelty scores from a wide range of linguistic features. They additionally report moderate inter-annotator agreement, underscoring that novelty itself is challenging to measure reliably by humans.
@@
The negative effect of model size observed in corpus-based datasets mirrors the inverse scaling effect reported in reading time studies oh_why_2023; WILCOX2025104650. Recently, oh-etal-2024-frequency argued that this effect is largely driven by word frequency, with larger models assigning increasingly non-human-like expectations to rare words. Since corpus-based novelty scores are often correlated with word frequency do-dinh-etal-2018-weeding; reimann_when_2024, frequency can be an underlying factor for this negative scaling effect.
By contrast, the synthetic datasets control for lexical identity and frequency: conventional and novel senses are elicited for the same set of words, and stimuli are constructed to cleanly separate the two senses. Under these controls, scaling improves alignment with novelty labels, agreeing with reports of positive scaling effects in other behavioural settings such as acceptability judgments and other reading time studies tjuatja-etal-2025-goes; wilcox-etal-2023-language.
@@
Our cloze-surprisal approach improves correlation across many model variants. In Figures 3 and 4, cloze-surprisal is found to be shifting the surprisal values away from near-zero values. In examples from Table 8, cloze-surprisal often raises the surprisal of the metaphorical word, consistent with the intuition that the full context is needed to process the metaphor. Moreover, in cases where the metaphorical word begins a sentence (e.g. example 14), direct surprisal is naturally high—regardless of novelty—whereas cloze-surprisal successfully moderates such inflated values.
Instruction-tuning—despite its goal of aligning model outputs with human intent—does not enhance the human-likeness of surprisal. In fact, similar to prior findings kuribayashi-etal-2024-psychometric, instruction-tuned models tend to reduce alignment between predicted probabilities and human annotation scores.
@@ 7 Conclusion
We acknowledge several limitations in this work. First, the scarcity of high-quality metaphor novelty annotations in existing literature constrains both coverage and generalizability. Second, we rely on pretrained language models whose training data and processes are not fully disclosed. Finally, although we carefully analyse the effects of model architecture, size, and domain (genre), future work could adopt mixed-effects models to test the interaction of these variables.
@@ A.1 Statistics and Examples
In Table 6, we describe in numbers the three datasets under study. In Table 8, we list 16 examples from the VUA-METANOV dataset, 6 examples from the LAI2009-METANOV dataset, and 4 examples from the GPT4o-METANOV dataset.
@@
VUA-ratings
Class Conventional Novel Score Surprisal Cloze
@@
Fiction 1. ‘ Tell him I am very sorry, but I must fill the quota. ’ 2. Adam might have escaped the file memories for years, suppressed them and jerked violently <14> by those events. c: -.441
n: .531
9.211
13.26
6.887
14.74
@@
3. It was an excitement that <11> and I had long dreamed of that scatter of tiny, magically named islands strewn across one third of a globe. 4. The seemingly random and <11> designed to disguise a boat’s shape from the prying eyes of U-Boat captains, so it <10> in the Bahamas. c: .278
n: .588
10.99
7.426
15.17
9.109
@@
News 5. One Mr Clarke can not duck away from if he wants to avoid a second Winter of Discontent .. 6. This was conveniently encapsulated in the first try. c: -.094
n: .500
3.969
8.586
7.344
14.35
@@
7. Thrusts of resistance ( mass demonstrations, resignations, tax rebellions, etc ) would come in crests. 8. Travel: A pilgrimage sans progress Elisabeth de Stroumillo potters round Poitou c: .382
n: .514
10.97
8.563
13.66
15.31
@@
Academic 9. The Tehuana dress is by no means the most decorative variant or the closest to pre-Hispanic forms of clothing. 10. Interwoven with these images are subtler references to the metaphorical borderlines which separate Latin American <5> and North America. c: -.194
n: .529
6.457
9.117
11.45
12.78
@@
11. This is often linked with a supposed denunciatory effect — the idea that the mandatory life sentence denounces murder as emphatically as possible <18> this crime. 12. He certainly held deep convictions as to the <9>, but at least a part of his apparent hostility was assumed for the occasion, a hard <7> in the end. c: .294
n: .514
9.961
5.648
8.109
9.469
@@
Conversat. 13. Me dad said he’s had enough Well, we were debating whether to give it to you or not. 14. Struggled with it a little c: -.633
n: .552
3.934
17.02
7.867
12.31
15. That’s an old trick. 16. Can you sort erm, madame out? c: .310
n: .567
4.063
8.734
12.34
10.13
LAI2009-METANOV
sank (V) 17. Upon hearing the news my spirits sank 18. Upon having the data my prediction sank 3.900
10.10
battle (N) 19. Those chess players are prepared for battle 20. Those plastic surgeons are prepared for battle 4.238
5.758
fragile (Adj) 21. His mental condition remains fragile 22. His website popularity remains fragile 6.508
9.023
GPT4o-METANOV
@@
anchor (N) 23. Her family was her emotional anchor during the crisis. 24. The smell of coffee became an anchor to mornings that no longer came. 3.615
7.488
@@
navigate (V) 25. The software helps users navigate complex legal documents. 26. He tried to navigate the silence like a sailor without stars. 3.941
8.391
@@
Table 8: Examples from each dataset. The metaphor word is in boldface within texts. For simpler presentations, we remove some words from long sentences and replace them with a tag of the number of words removed, e.g. <11>. Score is the novelty score, Surprisal is the GPT2-Base recorded surprisal, and Cloze is the GPT2-Base cloze-surprisal. In those three columns, values on top correspond to the conventional example, and on bottom to the novel example.
A.2 GPT4o-METANOV Construction

### v2

@@ 4.2 Evaluating the correlation
To determine whether LMs surprisal correlates with different metaphor novelty annotations, we measure multiple correlation metrics between surprisal and metaphor novelty scores/labels. For continuous novelty scores (as in VUA-ratings), we compute Pearson’s $r$ Pearson (1895) and Spearman’s $\rho$ Spearman (1904) correlations. For binary labels (conventional vs. novel), we compute the Rank-biserial $r_{b}$ Glass (1966) correlation. Rank-biserial estimates the probability that a random observation from the set of novel metaphors has a larger surprisal than one from the set of conventional metaphors, minus the reverse probability. We also estimate the potential of surprisal as a discriminator for binary novelty labels using the Area Under the ROC Curve (AUC) Fawcett (2006). We also compute the significance of all these estimates. All our metrics and tests do not assume normality, except for Pearson’s, for which we also provide a non-parametric alternative (Spearman).
@@ 4.3 Settings
We examine three families of decoder-only causal LMs (GPT-2, Llama 3 and Qwen2.5) Radford et al. (2019); Grattafiori et al. (2024); Yang et al. (2024). For each model family, we select 3-4 different sizes, to represent the effect of model size on the correlation. To investigate the effect of instruction-tuning on the correlation, we include the instruction-tuned variants of Llama 3 and Qwen2.5.
@@
As explained in Section 3, we perform experiments on four datasets (Tables 6 and 7). As VUA-ratings has continuous novelty scores, we evaluate the correlation of its continuous scores to surprisal. In addition, we convert these scores to binary labels using a threshold of $0.5$ and evaluate their correlation to surprisal, allowing us to compare the results of VUA-ratings to VUA-dictionary, Lai2009 and GPT-4o-metaphors, which only have binary novelty annotations.
@@ 4.4 Summary
Our experiments rely on a large collection of surprisal measures for each metaphor word in the four datasets: 32 different surprisal values (including direct and cloze) for 15,155 metaphoric words of the VUA dataset and for 208 and 200 metaphoric words of Lai2009 and GPT-4o-metaphors, respectively. In Figure 3, we plot the distributions of surprisal values by model, in addition to metaphor novelty scores/labels from VUA. From these plots and further normal distribution tests, we find that novelty scores and surprisal values are not strictly normally distributed.
@@ 5.1 Overall Correlation Results
Generally, we find LMs’ direct surprisal values correlate positively with metaphor novelty annotations across the four datasets (Tables 1 and 2). All correlation estimates are statistically significant for both the corpus-based and synthetic datasets  [^7] The relatively large corpus-based datasets (15,155 datapoints) can bias the significance values; however, we get similar significance values for the relatively small synthetic datasets (100 datapoints).. Overall for direct surprisal, the largest Pearson correlation $r=.419$, largest Spearman correlation $\rho=.417$, largest rank-biserial correlation $r_{b}=.638$ and largest $AUC=.819$ come from GPT2-base on VUA-ratings.
We also find that the correlation strengths’ scales differ across the different annotation datasets. By comparing the rank-biserial estimate $r_{b}$ across the four datasets, we find its ranges as follows: (.47-.64) for VUA-ratings, (.43-.58) for VUA-dictionary, (.28-.50) for Lai2009 and (.38-.63) for GPT-4o-metaphors. These results show that surprisal correlates relatively more with the corpus-based datasets than the synthetic datasets. Also, surprisal correlates relatively more with human ratings on VUA than the dictionary-based annotation approach. While the most controlled dataset (Lai2009) gets the weakest correlations with surprisal.
@@
Table 3: Instruction-tuning % gains (over corresponding base variants) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT-4o-metaphors.
@@ 5.2 Model Size Effects
The effect of model size is consistent (but diverging in direction) across the two dataset types. We plot the rank-biserial correlations of the ten model variants for the four datasets in Figure 2. The two corpus-based datasets (blue) show a monotonically decreasing correlation as model size increases per model family. On the other hand, the two synthetic datasets (red) exhibit an increasing correlation behaviour as model size increases, with a single minor exception at the GPT2 model family on the GPT-4o-metaphors dataset.
@@ 5.4 Cloze-surprisal
In Table 4, we report the percentage gains in rank-biserial correlations for the cloze-surprisal values over the direct-surprisal values from the same model variants across the four datasets. We find cloze-surprisal to be boosting the correlations with extra points in many cases; however, it also deteriorates the correlations heavily in other cases. In Figure 3, we can observe the distribution of cloze-surprisal of the metaphoric words shifted away from zero and approaching a normal distribution. Also in Figure 4, we observe the same effect. However, against our intuition, the effect of cloze-surprisal is not more impactful in cases where the metaphor is located earlier in the sentence than later. The effect of cloze-surprisal, although positive, is not related to the position of the metaphoric word in the sentence.
For the GPT2 model family, cloze-surprisal consistently improves the corpus-based datasets’ correlations, achieving the strongest correlations in this study (out of the 32 recorded surprisals on each dataset) coming from GPT2-base on VUA-ratings with values of $r=.499$, $\rho=.499$, $r_{b}=.687$ and $auc=.843$. However, it deteriorates the correlations at large scales on the synthetic datasets.
Llama and Qwen model families also show positive gains in most of the cases. Most importantly, unlike the GPT2 model family, it boosts correlations on the synthetic datasets except for a few cases. Cloze-surprisal achieves the overall strongest correlation on the Lai2009 dataset using the Qwen-14B model, with $r_{b}=.539$ and $auc=.779$. Also, it achieves the overall strongest correlation on the GPT-4o-metaphors dataset using the Llama-8B model, with $r_{b}=.684$ and $auc=.842$.
Model VUA-r VUA-d Lai GPT-4o
@@
Table 4: Cloze-surprisal % gains (over direct-surprisal from the same model) in Rank-biserial’s correlation estimates between surprisal and novelty scores/labels in the four datasets: VUA-ratings, VUA-dictionary, Lai2009 and GPT-4o-metaphors.
@@ 6 Discussion
Across multiple novelty annotation setups—continuous human ratings, dictionary-based binary labels, linguist-designed contrasts, and LLM-generated conventional vs. novel senses—we find consistently moderate associations between LM surprisal and metaphor novelty (best: $r=.49,\rho=.50,r_{b}=.69,\mathrm{AUC}=.84$). While these results are not directly comparable to prior work due to differences in datasets and task formulations, their magnitudes are broadly in line with reported surprisal–behavior associations in acceptability judgments Tjuatja et al. (2025) and reading times Oh and Schuler (2023a). The closest point of reference in the metaphor novelty literature is Parde and Nielsen (2018), who report $r=.44$ when predicting novelty scores from a wide range of linguistic features. They additionally report moderate inter-annotator agreement, underscoring that novelty itself is challenging to measure reliably by humans.
@@
The negative effect of model size observed in corpus-based datasets mirrors the inverse scaling effect reported in reading time studies Oh and Schuler (2023b); Wilcox et al. (2025). Recently, Oh et al. (2024) argued that this effect is largely driven by word frequency, with larger models assigning increasingly non-human-like expectations to rare words. Since corpus-based novelty scores are often correlated with word frequency Do Dinh et al. (2018); Reimann and Scheffler (2024), frequency can be an underlying factor for this negative scaling effect.
By contrast, the synthetic datasets control for lexical identity and frequency: conventional and novel senses are elicited for the same set of words, and stimuli are constructed to cleanly separate the two senses. Under these controls, scaling improves alignment with novelty labels, agreeing with reports of positive scaling effects in other behavioural settings such as acceptability judgments and other reading time studies Tjuatja et al. (2025); Wilcox et al. (2023).
@@
Our cloze-surprisal approach improves correlation across many model variants. In Figures 3 and 4, cloze-surprisal is found to be shifting the surprisal values away from near-zero values and pushing towards a normal distribution.
In examples from Table 8, cloze-surprisal often raises the surprisal of the metaphorical word, consistent with the intuition that the full context is needed to process the metaphor and realise its true predictability. Moreover, in cases where the metaphorical word begins a sentence (e.g. example 14), direct surprisal is naturally high—regardless of novelty—whereas cloze-surprisal successfully moderates such inflated values.
Instruction-tuning—despite its goal of aligning model outputs with human intent—does not enhance the human-likeness of surprisal. In fact, similar to prior findings Kuribayashi et al. (2024), instruction-tuned models tend to reduce alignment between predicted probabilities and human annotation scores.
@@ 7 Conclusion
We acknowledge several limitations in this work. First, the scarcity of high-quality metaphor novelty annotations in existing literature constrains both coverage and generalizability. Second, we rely on pretrained language models whose training data and processes are not fully disclosed.
Finally, although we carefully analyse the effects of model architecture, size, and domain (genre), future work could adopt mixed-effects models to test the interaction of these variables.
Acknowledgments
This research has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – CRC-1646, project number 512393437, project A05.
We acknowledge the reviewers for their valuable comments and feedback. And we also thank Vicky Tzuyin Lai and Tiago Pimentel for their responsiveness to our inquiries.
@@ A.1 Statistics and Examples
In Tables 6 and 7, we describe in numbers the four datasets under study. In Table 8, we list 16 examples from the VUA datasets, 6 examples from the LAI2009 dataset, and 4 examples from the GPT-4o-metaphors dataset.
@@
VUA datasets
Sentence ratings dictionary direct cloze
1. ‘ Tell him I am very sorry, but I must fill the quota. ’ -0.441 conventional 9.043 7.317
@@
2. Adam might have escaped the file memories for years, suppressed them and jerked violently <14> by those events. 0.531 novel 14.23 14.69
@@
3. It was an excitement that <11> and I had long dreamed of that scatter of tiny, magically named islands strewn across one third of a globe. 0.278 conventional 10.17 16.08
4. The seemingly random and <11> designed to disguise a boat’s shape from the prying eyes of U-Boat captains, so it <10> in the Bahamas. 0.588 conventional 7.903 12.46
@@
5. One Mr Clarke can not duck away from if he wants to avoid a second Winter of Discontent .. -0.094 conventional 3.165 7.454
6. This was conveniently encapsulated in the first try. 0.500 conventional 7.918 14.54
@@
7. Thrusts of resistance ( mass demonstrations, resignations, tax rebellions, etc ) would come in crests. 0.382 novel 12.46 16.84
8. Travel: A pilgrimage sans progress Elisabeth de Stroumillo potters round Poitou 0.514 conventional 8.188 14.81
9. The Tehuana dress is by no means the most decorative variant or the closest to pre-Hispanic forms of clothing. -0.194 conventional 6.466 11.92
@@
10. Interwoven with these images are subtler references to the metaphorical borderlines which separate Latin American <5> and North America. 0.529 conventional 10.16 12.57
@@
11. This is often linked with a supposed denunciatory effect — the idea that the mandatory life sentence denounces murder as emphatically as possible <18> this crime. 0.294 conventional 11.02 12.61
12. He certainly held deep convictions as to the <9>, but at least a part of his apparent hostility was assumed for the occasion, a hard <7> in the end. 0.514 conventional 5.662 9.781
@@
13. Me dad said he’s had enough Well, we were debating whether to give it to you or not. -0.633 conventional 3.752 8.038
14. Struggled with it a little 0.552 conventional 17.13 14.92
15. That’s an old trick. 0.310 conventional 4.013 11.92
16. Can you sort erm, madame out? 0.567 conventional 8.820 9.591
LAI2009
Sentence label direct cloze
17. Upon hearing the news my spirits sank conventional 4.358 11.82
18. Upon having the data my prediction sank novel 10.28 13.27
19. Those chess players are prepared for battle conventional 5.373 12.32
20. Those plastic surgeons are prepared for battle novel 6.691 13.28
21. His mental condition remains fragile conventional 7.251 13.34
22. His website popularity remains fragile novel 9.538 13.95
GPT-4o-metaphors
23. Her family was her emotional anchor during the crisis. conventional 4.322 14.12
@@
24. The smell of coffee became an anchor to mornings that no longer came. novel 7.518 12.42
@@
25. The software helps users navigate complex legal documents. conventional 3.830 5.575
26. He tried to navigate the silence like a sailor without stars. novel 8.041 8.402
@@
Table 8: Examples from each dataset. The metaphor word is in boldface within sentences. For simpler presentations, we remove some words from long sentences and replace them with a tag of the number of words removed, e.g. <11>. rating is the Do Dinh et al. (2018) score, and dictionary is the Reimann and Scheffler (2024) annotation. direct is the GPT2-Base recorded surprisal, and cloze is the GPT2-Base cloze-surprisal. Paired examples are picked randomly for VUA datasets from the same genres to illustrate the differences between conventional and novel instances according to “ratings”. Sentences 1-4 are from Fiction, 5-8 from News, 9-12 from Academic and 13-16 from Conversation. Also obviously, Do Dinh et al. (2018) ratings and Reimann and Scheffler (2024) annotations do not agree in many cases.
A.2 GPT-4o-metaphors Construction